# Balancing Quality and Variation: Spam Filtering Distorts Data Label Distributions

### v1

@@ Abstract
For datasets to accurately represent diverse opinions in a population, they must preserve variation in data labels while filtering out spam or low-quality responses. How can we balance annotator reliability and representation? We empirically evaluate how a range of heuristics for annotator filtering affect the preservation of variation on subjective tasks. We find that these methods, designed for contexts in which variation from a single ground-truth label is considered noise, often remove annotators who disagree instead of spam annotators, introducing suboptimal tradeoffs between accuracy and label diversity. We find that conservative settings for annotator removal (<5%) are best, after which all tested methods increase the mean absolute error from the true average label. We analyze performance on synthetic spam to observe that these methods often assume spam annotators are less random than real spammers tend to be: most spammers are distributionally indistinguishable from real annotators, and the minority that are distinguishable tend to give fixed answers, not random ones. Thus, tasks requiring the preservation of variation reverse the intuition of existing spam filtering methods: spammers tend to be less random than non-spammers, so metrics that assume variation is spam fare worse. These results highlight the need for spam removal methods that account for label diversity.
@@ 1 Introduction
We examine the effects of applying several common methods for discounting spam annotators based on their labeling behavior. We test three of these methods—MACE Hovy et al. (2013), CrowdTruth Aroyo and Welty (2014), and inter-annotator agreement metrics—on relatively subjective tasks and analyze effects on variability in the filtered data. We find that, although many methods are near-indistinguishable in terms of their accuracy at classifying spam annotators, some are far more likely to remove non-spam annotators who disagree. Furthermore, we find that under most tested methods, removing more annotators degrades the variety of opinions expressed, without improving accuracy at removing spam annotators; thus, these methods seem most effective only when a very low number of annotators are removed.
@@
Despite the existence of spam removal methods that use attention checks or metadata (e.g., time spent on task), filtering based on labeling behavior remains common practice Klie et al. (2024); thus, weaknesses in these methods risk affecting a wide range of common machine learning tasks. Our results indicate that spam detection for subjective problems flips the assumptions of existing methods: spam annotators are often less random than non-spam ones. Thus, attempts to remove spam can backfire by instead removing annotators with minority opinions who are not spammers. As a result, existing methods work best when only low percentages of annotators are removed based on their labeling behavior. When over-filtering for spam, these methods risk distorting the distributions of labels.
@@ 2 Related Work
Drawing a conceptual boundary between spammers and genuine annotators is complex. Definitions vary on what range of intentional, inattentive, or low-effort behaviors should be filtered out, and on whether spammers are posited as too random or too fixed. Buchholz and Latorre (2011) highlight that spammers are incentivized to earn more money faster, leading them to ignore task instructions or participation requirements. Rothwell et al. (2015) argue that spammers act with intention, unlike other types of low-quality annotators, and show repeated patterns in an attempt to complete tasks fast. In contrast, Raykar and Yu (2012) posit that spammers assign labels randomly, because they do not follow labeling criteria, skip reading the instances or might use automation. Gadiraju et al. (2015) present a nuanced taxonomy of annotator types and underscore that genuine annotators’ behavior might overlap with spammers, e.g., failing attention checks for innocuous reasons. The datasets used in our study excluded annotators if they failed data quality checks combining multiple sources of information, thus following a wider definition of spam (Aroyo et al., 2023; Huang et al., 2023, see Section 3). Any definition of “spammer” includes or excludes different subsets of annotators; these ambiguous boundaries suggest that different subsets of spammers may exhibit different behaviors, potentially raising challenges in distinguishing spammers from non-spammers.
@@
Data quality and questionable trust in non-expert raters are longstanding problems in crowdsourced annotation Snow et al. (2008). Attempts to improve data quality may modify tasks to attract less spam before data collection Eickhoff and de Vries (2013) or use quality control afterwards Difallah et al. (2012). Methods for a posteriori detection of low-quality raters and spammers often use intrinsic metrics based on the labeling behavior itself Buchholz and Latorre (2011). Intrinsic metrics used for spammer detection include clustering on a post-processed annotation matrix Traganitis and Giannakis (2021), rater similarity and agreement scores Ak et al. (2021), or analyzing sequential spamming behaviors Ba et al. (2024), among others Ipeirotis et al. (2010); Raykar and Yu (2012); Gadiraju et al. (2015). Other methods analyze labeling behavior with the goal of aggregating to the true label while accounting for varying annotator reliability. Dawid and Skene (1979) model annotator error rates to estimate the true labels and are foundational to many subsequent aggregation methods Whitehill et al. (2009); Welinder et al. (2010), including in NLP Wiebe et al. (1999). Passonneau and Carpenter (2014) present a probabilistic variant of the Dawid & Skene model, and many other extensions of this basic model exist Paun et al. (2018, 2022). In particular, Hovy et al. (2013) present MACE, a probabilistic model tailored towards estimating annotator competence by modeling spamming behaviors. In contrast, CrowdTruth, a non-probabilistic paradigm, derives quality metrics from vector space representations of annotators, annotated examples and annotations Aroyo and Welty (2014); Dumitrache et al. (2018b). We evaluate MACE and CrowdTruth as they underwent widespread adoption in NLP and have reference implementations available (see Sections 4.1 and 4.2).
@@
There is a growing body of work researching informative disagreement, diversity of perspectives, and label variation in human annotation (Plank, 2022; Leonardelli et al., 2023; Sandri et al., 2023; Frenda et al., 2024; Fleisig et al., 2024). These studies agree that aggregating labels into a single truth is an oversimplification for many tasks (Aroyo and Welty, 2015; Uma et al., 2021; Basile et al., 2021) and might not represent perspectives fairly (Abercrombie et al., 2022). Instead, studies release annotator-level labels Prabhakaran et al. (2021) to enable alternative approaches, such as modeling individual annotators’ rating behaviors Fleisig et al. (2023); Orlikowski et al. (2023); Heinisch et al. (2023); Orlikowski et al. (2025). Our work is motivated by studies on rating distributions in a given population as an alternative to single ground truth prediction Sorensen et al. (2024); Meister et al. (2025). Among these, Prabhakaran et al. (2024) study systematic disagreement using similar metrics to ours, but on the level of demographic subgroups. In this context, the issue of how capturing labeling variation intersects with annotation quality is largely unexplored. One exception is VariErr Weber-Genzel et al. (2024), an annotation methodology to differentiate between annotation errors and plausible variation in annotation. In contrast, we study properties of methods that determine annotator reliability, not individual annotation errors.
@@ 3 Datasets
We selected two datasets for the basis of our experiments: DICES 350 Aroyo et al. (2023) and Huang et al. (2023)’s survey of Amazon Mechanical Turk workers. We present each dataset’s statistics and discuss our dataset selection process below.
@@
DICES-350 Aroyo et al. (2023), a harmful language dataset, consists of 43,050 annotations on a 3-point scale across 350 items, each of which was labeled by every participant. 123 annotators participated, of whom 19 annotators were labeled as spam (15% of annotators).
@@
From Huang et al. (2023)’s survey, we used 16 questions on a 7-point scale, each of which was answered by every participant, for a total of 3,312 annotations. 207 annotators participated, of whom 40 were labeled as spam (19% of annotators).
@@
Our experiments require datasets that retain (a) responses from multiple annotators per question, permitting measurement of disagreement statistics, and (b) responses from known spammers. Despite increasing availability of annotator-level data, [^1] For example, https://github.com/mainlp/awesome-human-label-variation most public datasets do not include spammer information. For papers that report spammer removal, we contacted authors for access to the unfiltered datasets, but spammer responses are regularly lost over time (e.g., Buchholz and Latorre, 2011; Dumitrache et al., 2017; Paun et al., 2018). Even for published data, maintaining data access is not always possible; some datasets with verified spammer information were no longer available (e.g., Soberón et al., 2013; Gadiraju et al., 2015). Similarly, many studies on spammer detection evaluate only on downstream performance or exclusively use synthetic data, so they do not provide metadata on known natural spammers (e.g., Raykar and Yu, 2012; Ak et al., 2021). See Appendix D for details on all 22 considered datasets, including spammer metadata and data availability. In summary, DICES-350 and the MTurk survey are, to our knowledge, the only available datasets meeting our criteria. Nevertheless, these datasets do represent two representative use cases in which preserving rater variation is essential. DICES-350 collects annotations on AI safety to preserve variation on diverse perspectives regarding high-stakes topics; the MTurk dataset polls workers on their personal opinions about their crowdwork experiences in order to understand the range of opinions of the community.
@@ 4.1 Multi-Annotator Competence Estimation (MACE)
MACE Hovy et al. (2013) is based on a probabilistic model of annotation. We highlight a few aspects of MACE that are important to our study and refer to the original paper for full details. The model includes a parameter $\theta$ for each annotator which encodes the probability that they give the true answer (competence). Specifically, for each instance $i$ and annotator $j$, the binary variable $S_{ij}$ indicates whether an annotator is spamming. $S_{ij}$ is drawn from a Bernoulli distribution with parameter $1-\theta_{j}$. If the annotator is spamming, i.e., $S_{ij}=1$, then the assigned label $A_{ij}$ is sampled from a multinomial distribution with a parameter vector $\zeta_{j}$ that encodes each annotator’s spamming strategy. Otherwise, if $S_{ij}=0$, the model assumes that the annotator simply assigns the correct label—an intentional simplification to focus on modeling spam behavior. Only the annotations $A_{ij}$ are observed; the other parameters are inferred when updating the model from data.
Usually, when applying MACE for label aggregation, the model would weigh all annotations to estimate the correct labels without discarding specific annotators. But the learned parameters can also be used to identify spamming annotators: the competence $\theta$ correlates more strongly than agreement measures with an annotator’s fraction of correctly annotated examples Hovy et al. (2013) and both learned annotator parameters ($\theta$, $\zeta$) were shown to encode characteristic spamming behaviors Paun et al. (2018). Consequently, other studies have used MACE to exclude spammers during dataset construction based on an empirically chosen threshold for competence Pei and Jurgens (2023). In our experiments, we also use the competence parameter to score annotators.
@@ 4.2 CrowdTruth
The CrowdTruth framework Aroyo and Welty (2014) computes several interdependent quality metrics that use vector representations of annotations to measure disagreement and ambiguity, including a worker quality score. The metrics follow the aim of ambiguity-aware label aggregation, so that, for example, disagreement on ambiguous instances discounts worker quality less. The worker quality score (WQS) for an annotator $i$ is computed as the product of two other scores $WQS(i)=WUA(i)\cdot WWA(i)$, the worker-unit agreement (WUA) and the worker-worker agreement (WWA). Conceptually, WWA measures how similar a given worker’s annotations are to other workers, weighted by the workers’ quality and the instances’ ambiguity. WUA measures how much a worker agrees with the aggregate label over all their annotated instances, weighted by the instances’ ambiguity. (See Appendix A for details on how these metrics are computed.)
The CrowdTruth metrics were explored on various tasks Dumitrache et al. (2017, 2018a) and have been explicitly used for spammer removal Dumitrache et al. (2021). In a related study, Soberón et al. (2013) report an accuracy of $0.88$ for removing spam annotators using CrowdTruth metrics. In our experiments, we use the worker quality score (WQS) to score annotators.
@@ 4.3 Cohen’s Kappa
As a representative example of using inter-annotator agreement metrics to filter annotators, we compute each annotator’s pair-wise agreement as measured by Cohen’s kappa Cohen (1968) with each other annotator. We then use the averaged agreement to score annotators.
@@ 6 Discussion and Conclusion
Since different types of low-quality annotators behave differently, annotators that need to be excluded can exhibit varied behaviors beyond simple patterns such as always selecting the same answer. Consequently, while annotator reliability scoring can often single out spammers showing these stereotypical behaviors, many genuine annotators will be scored similarly to low-quality raters. This result highlights that in addition to the labeling behavior, additional signals should be included in spammer removal. These can be metadata, such as when and how much time is spent on annotation Rothwell et al. (2015) or previous acceptance rates of annotators Difallah et al. (2012). Similarly, verifiable test questions could be used, that is, unambiguous cases where comparison to known answers is possible (gold standard or attention checks, Difallah et al., 2012; Rothwell et al., 2015).
@@
In addition, the scarcity of available metadata on removed spam data makes it difficult to characterize spammer behavior across a range of contexts. Difallah et al. (2012) highlight a “need for new benchmarks on which to evaluate and compare existing and novel spam detection techniques for crowdsourcing platforms” that still persists. Datasets often do not report spam filtering techniques or preserve the spam responses; however, this data is extremely helpful for more fine-grained characterization of spam behavior, especially in complex contexts where variation is expected. Thus, making this data available would be a valuable resource for future research.
@@
Categorizing raters as “spammers” is based on varying definitions and procedures: “gold spammers” are not ground truth the same way that other data might be. Importantly, self-reported spammer information, where spammers disclose themselves, is largely not even gathered (for an exception, see Paun et al., 2018) and not publicly available. Consequently, the “gold spammer” labels used in our study are based on external categorizations. While these are reported to be based on manual checks and multiple data types (labeling behavior, metadata, and attention checks), there remains a risk of wrong categorizations.

### v2

@@ Abstract
For datasets to accurately represent diverse opinions in a population, they must preserve variation in data labels while filtering out spam or low-quality responses. How can we balance annotator reliability and representation? We empirically evaluate how a range of heuristics for annotator filtering affect the preservation of variation on subjective tasks. We find that these methods, designed for contexts in which variation from a single ground-truth label is considered noise, often remove annotators who disagree instead of spam annotators, introducing suboptimal tradeoffs between accuracy and label diversity. We find that conservative settings for annotator removal (<5%) are best, after which all tested methods increase the mean absolute error from the true average label. We analyze performance on synthetic spam to observe that these methods often assume spam annotators are more random than real spammers tend to be: most spammers are distributionally indistinguishable from real annotators, and the minority that are distinguishable tend to give relatively fixed answers, not random ones. Thus, tasks requiring the preservation of variation reverse the intuition of existing spam filtering methods: spammers tend to be less random than non-spammers, so metrics that assume variation is spam fare worse. These results highlight the need for spam removal methods that account for label diversity.
@@ 1 Introduction
We examine the effects of applying several common methods for discounting spam annotators based on their labeling behavior. We test three of these methods—MACE hovy-etal-2013-learning, CrowdTruth aroyo_three_2014, and inter-annotator agreement metrics—on relatively subjective tasks and analyze effects on variability in the filtered data. We find that, although many methods are near-indistinguishable in terms of their accuracy at classifying spam annotators, some are far more likely to remove non-spam annotators who disagree. Furthermore, we find that under most tested methods, removing more annotators degrades the variety of opinions expressed, without improving accuracy at removing spam annotators; thus, these methods seem most effective only when a very low number of annotators are removed.
@@
Despite the existence of spam removal methods that use attention checks or metadata (e.g., time spent on task), filtering based on labeling behavior remains common practice klie-etal-2024-analyzing; thus, weaknesses in these methods risk affecting a wide range of common machine learning tasks. Our results indicate that spam detection for subjective problems flips the assumptions of existing methods: spam annotators are often less random than non-spam ones. Thus, attempts to remove spam can backfire by instead removing annotators with minority opinions who are not spammers. As a result, existing methods work best when only low percentages of annotators are removed based on their labeling behavior. When over-filtering for spam, these methods risk distorting the distributions of labels.
@@ 2 Related Work
Drawing a conceptual boundary between spammers and genuine annotators is complex. Definitions vary on what range of intentional, inattentive, or low-effort behaviors should be filtered out, and on whether spammers are posited as too random or too fixed. buchholz_crowdsourcing_2011 highlight that spammers are incentivized to earn more money faster, leading them to ignore task instructions or participation requirements. rothwell_controlling_2015 argue that spammers act with intention, unlike other types of low-quality annotators, and show repeated patterns in an attempt to complete tasks fast. In contrast, raykar_eliminating_2012 posit that spammers assign labels randomly, because they do not follow labeling criteria, skip reading the instances or might use automation. gadiraju_understanding_2015 present a nuanced taxonomy of annotator types and underscore that genuine annotators’ behavior might overlap with spammers, e.g., failing attention checks for innocuous reasons. The datasets used in our study excluded annotators if they failed data quality checks combining multiple sources of information, thus following a wider definition of spam (aroyo2023dices; huang-etal-2023-incorporating, see Section 3). Any definition of “spammer” includes or excludes different subsets of annotators; these ambiguous boundaries suggest that different subsets of spammers may exhibit different behaviors, potentially raising challenges in distinguishing spammers from non-spammers.
@@
Data quality and questionable trust in non-expert raters are longstanding problems in crowdsourced annotation snow-etal-2008-cheap. Attempts to improve data quality may modify tasks to attract less spam before data collection eickhoff_increasing_2013 or use quality control afterwards difallah_etal_2012_mechanical_cheat. Methods for a posteriori detection of low-quality raters and spammers often use intrinsic metrics based on the labeling behavior itself buchholz_crowdsourcing_2011. Intrinsic metrics used for spammer detection include clustering on a post-processed annotation matrix traganitis_identifying_2021, rater similarity and agreement scores ak_spammer_2021, or analyzing sequential spamming behaviors ba_data_2024, among others ipeirotis_quality_2010; raykar_eliminating_2012; gadiraju_understanding_2015. Other methods analyze labeling behavior with the goal of aggregating to the true label while accounting for varying annotator reliability. dawid_maximum_1979 model annotator error rates to estimate the true labels and are foundational to many subsequent aggregation methods whitehill_whose_2009; welinder_multidimensional_2010, including in NLP wiebe-etal-1999-development. passonneau-carpenter-2014-benefits present a probabilistic variant of the Dawid & Skene model, and many other extensions of this basic model exist paun-etal-2018-comparing; paun_statistical_2022. In particular, hovy-etal-2013-learning present MACE, a probabilistic model tailored towards estimating annotator competence by modeling spamming behaviors. In contrast, CrowdTruth, a non-probabilistic paradigm, derives quality metrics from vector space representations of annotators, annotated examples and annotations aroyo_three_2014; CrowdTruth2. We evaluate MACE and CrowdTruth as they underwent widespread adoption in NLP and have reference implementations available (see Sections 4.1 and 4.2).
@@
There is a growing body of work researching informative disagreement, diversity of perspectives, and label variation in human annotation (plank-2022-problem; leonardelli-etal-2023-semeval; sandri-etal-2023-dont; frenda_perspectivist_2024; fleisig-etal-2024-perspectivist). These studies agree that aggregating labels into a single truth is an oversimplification for many tasks (aroyo_truth_2015; uma-2021-survey; basile-etal-2021-need) and might not represent perspectives fairly (nlperspectives-2022-perspectivist). Instead, studies release annotator-level labels prabhakaran-etal-2021-releasing to enable alternative approaches, such as modeling individual annotators’ rating behaviors fleisig-etal-2023-majority; orlikowski-etal-2023-ecological; heinisch-etal-2023-architectural; orlikowski-etal-2025-beyond. Our work is motivated by studies on rating distributions in a given population as an alternative to single ground truth prediction sorensen_etal_2025_pluralistic; meister-etal-2025-benchmarking. Among these, prabhakaran-etal-2024-grasp study systematic disagreement using similar metrics to ours, but on the level of demographic subgroups. In this context, the issue of how capturing labeling variation intersects with annotation quality is largely unexplored. One exception is VariErr weber-genzel-etal-2024-varierr, an annotation methodology to differentiate between annotation errors and plausible variation in annotation. In contrast, we study properties of methods that determine annotator reliability, not individual annotation errors.
@@ 3 Datasets
We selected two datasets for the basis of our experiments: DICES 350 aroyo2023dices and huang-etal-2023-incorporating’s survey of Amazon Mechanical Turk workers. We present each dataset’s statistics and discuss our dataset selection process below.
@@
DICES-350 aroyo2023dices, a harmful language dataset, consists of 43,050 annotations on a 3-point scale across 350 items, each of which was labeled by every participant. 123 annotators participated, of whom 19 annotators were labeled as spam (15% of annotators).
@@
From huang-etal-2023-incorporating’s survey, we used 16 questions on a 7-point scale, each of which was answered by every participant, for a total of 3,312 annotations. 207 annotators participated, of whom 40 were labeled as spam (19% of annotators).
@@
Our experiments require datasets that retain (a) responses from multiple annotators per question, permitting measurement of disagreement statistics, and (b) responses from known spammers. Despite increasing availability of annotator-level data, [^1] For example, https://github.com/mainlp/awesome-human-label-variation most public datasets do not include spammer information. For papers that report spammer removal, we contacted authors for access to the unfiltered datasets, but spammer responses are regularly lost over time (e.g., buchholz_crowdsourcing_2011; dumitrache_false_2017; paun-etal-2018-comparing). Even for published data, maintaining data access is not always possible; some datasets with verified spammer information were no longer available (e.g., soberon_measuring_2013; gadiraju_understanding_2015). Similarly, many studies on spammer detection evaluate only on downstream performance or exclusively use synthetic data, so they do not provide metadata on known natural spammers (e.g., raykar_eliminating_2012; ak_spammer_2021). See Appendix D for details on all 22 considered datasets, including spammer metadata and data availability. In summary, DICES-350 and the MTurk survey are, to our knowledge, the only available datasets meeting our criteria. Nevertheless, these datasets do represent two representative use cases in which preserving rater variation is essential. DICES-350 collects annotations on AI safety to preserve variation on diverse perspectives regarding high-stakes topics; the MTurk dataset polls workers on their personal opinions about their crowdwork experiences in order to understand the range of opinions of the community.
@@ 4.1 Multi-Annotator Competence Estimation (MACE)
MACE hovy-etal-2013-learning is based on a probabilistic model of annotation. We highlight a few aspects of MACE that are important to our study and refer to the original paper for full details. The model includes a parameter $\theta$ for each annotator which encodes the probability that they give the true answer (competence). Specifically, for each instance $i$ and annotator $j$, the binary variable $S_{ij}$ indicates whether an annotator is spamming. $S_{ij}$ is drawn from a Bernoulli distribution with parameter $1-\theta_{j}$. If the annotator is spamming, i.e., $S_{ij}=1$, then the assigned label $A_{ij}$ is sampled from a multinomial distribution with a parameter vector $\zeta_{j}$ that encodes each annotator’s spamming strategy. Otherwise, if $S_{ij}=0$, the model assumes that the annotator simply assigns the correct label—an intentional simplification to focus on modeling spam behavior. Only the annotations $A_{ij}$ are observed; the other parameters are inferred when updating the model from data.
Usually, when applying MACE for label aggregation, the model would weigh all annotations to estimate the correct labels without discarding specific annotators. But the learned parameters can also be used to identify spamming annotators: the competence $\theta$ correlates more strongly than agreement measures with an annotator’s fraction of correctly annotated examples hovy-etal-2013-learning and both learned annotator parameters ($\theta$, $\zeta$) were shown to encode characteristic spamming behaviors paun-etal-2018-comparing. Consequently, other studies have used MACE to exclude spammers during dataset construction based on an empirically chosen threshold for competence pei-jurgens-2023-annotator. In our experiments, we also use the competence parameter to score annotators.
@@ 4.2 CrowdTruth
The CrowdTruth framework aroyo_three_2014 computes several interdependent quality metrics that use vector representations of annotations to measure disagreement and ambiguity, including a worker quality score. The metrics follow the aim of ambiguity-aware label aggregation, so that, for example, disagreement on ambiguous instances discounts worker quality less. The worker quality score (WQS) for an annotator $i$ is computed as the product of two other scores $WQS(i)=WUA(i)\cdot WWA(i)$, the worker-unit agreement (WUA) and the worker-worker agreement (WWA). Conceptually, WWA measures how similar a given worker’s annotations are to other workers, weighted by the workers’ quality and the instances’ ambiguity. WUA measures how much a worker agrees with the aggregate label over all their annotated instances, weighted by the instances’ ambiguity. (See Appendix A for details on how these metrics are computed.)
The CrowdTruth metrics were explored on various tasks dumitrache_false_2017; dumitrache_crowdsourcing_2018 and have been explicitly used for spammer removal dumitrache_empirical_2021. In a related study, soberon_measuring_2013 report an accuracy of $0.88$ for removing spam annotators using CrowdTruth metrics. In our experiments, we use the worker quality score (WQS) to score annotators.
@@ 4.3 Cohen’s Kappa
As a representative example of using inter-annotator agreement metrics to filter annotators, we compute each annotator’s pair-wise agreement as measured by Cohen’s kappa cohen_weighted_1968 with each other annotator. We then use the averaged agreement to score annotators.
@@ 6 Discussion and Conclusion
Since different types of low-quality annotators behave differently, annotators that need to be excluded can exhibit varied behaviors beyond simple patterns such as always selecting the same answer. Consequently, while annotator reliability scoring can often single out spammers showing these stereotypical behaviors, many genuine annotators will be scored similarly to low-quality raters. This result highlights that in addition to the labeling behavior, additional signals should be included in spammer removal. These can be metadata, such as when and how much time is spent on annotation rothwell_controlling_2015 or previous acceptance rates of annotators difallah_etal_2012_mechanical_cheat. Similarly, verifiable test questions could be used, that is, unambiguous cases where comparison to known answers is possible (gold standard or attention checks, difallah_etal_2012_mechanical_cheat; rothwell_controlling_2015).
@@
In addition, the scarcity of available metadata on removed spam data makes it difficult to characterize spammer behavior across a range of contexts. difallah_etal_2012_mechanical_cheat highlight a “need for new benchmarks on which to evaluate and compare existing and novel spam detection techniques for crowdsourcing platforms” that still persists. Datasets often do not report spam filtering techniques or preserve the spam responses; however, this data is extremely helpful for more fine-grained characterization of spam behavior, especially in complex contexts where variation is expected. Thus, making this data available would be a valuable resource for future research.
@@
Categorizing raters as “spammers” is based on varying definitions and procedures: “gold spammers” are not ground truth the same way that other data might be. Importantly, self-reported spammer information, where spammers disclose themselves, is largely not even gathered (for an exception, see paun-etal-2018-comparing) and not publicly available. Consequently, the “gold spammer” labels used in our study are based on external categorizations. While these are reported to be based on manual checks and multiple data types (labeling behavior, metadata, and attention checks), there remains a risk of wrong categorizations.